OpenStudy (anonymous):

find the point on the line -6x+7y-5=0 which is closest to the point (1,3)

OpenStudy (anonymous):

the closest point between a line and a piont lies on a perpendicular line that connect the point and the line so for the line we should find the slope -4x + 7 y + 6 =0 solve for y 7y = 4x -6 y = 4/7 x - 6/7 so slope is m1= 4/7 now perpendicular lines have slopes that are negative reciprocals so m1= -1/m2 so 4/7 = - 1/ m2 m2= -7/4 now we have a slope m=-7/4 and a point (-5, 1) use slope point formula y-y, = m(x-x,) y- 1 = -7/4 (x+5) y-1 = -7/4 x - 35/4 y = -7/4 x - 35/4 +1 y= -7/4 X - 35/4 + 4/4 Y = -7/4 X - 31/4

OpenStudy (anonymous):

please give medal, if helped

OpenStudy (anonymous):

but what is the point?

OpenStudy (anonymous):

last answer on bottom

OpenStudy (anonymous):

I put that it but its asking me for a point not an equation

OpenStudy (anonymous):

Refer to the solution and plot from Mathematica 9, that is attached.